OpenStudy (anonymous):

3sqrt12 times sqrt 6

OpenStudy (anonymous):

what would the simplified version of 3 sqrt(12) be?

OpenStudy (anonymous):

3 sqrt 2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but ill tell you.... First you must leave out the 3 in the equation "3sqrt(12)", then you have to divide by the lowest number you can find. 12 / 2 = 6 / 2 = 3. but because there are 2 2's that means that you can simplify the equation "3sqrt(12)" into 6sqrt(3)

OpenStudy (anonymous):

so 6sqrt(3) * sqrt(6) is???

OpenStudy (anonymous):

18 sqrt 2